# AIR-Bench: Benchmarking Large Audio-Language Models via Generative Comprehension

### v1

@@ 4 Experiments
Table 4: The success rate of different strategies of matching hypotheses with the golden choices for the foundation benchmark.
@@ 4.1 Models
We evaluate the performance of various LALMs with instruction-following capabilities. These models are either open-sourced or accessible through public APIs, such as SpeechGPT (Zhang et al., 2023), BLSP (Wang et al., 2023a), SALMONN (Tang et al., 2023a), Qwen-Audio-Chat (Chu et al., 2023), and Qwen-Audio Turbo  [^3] https://help.aliyun.com/zh/dashscope/developer-reference/qwen-audio-api. Additionally, we consider large multi-modality models with audio understanding abilities like PandaGPT (Su et al., 2023), Macaw-LLM (Lyu et al., 2023), and NExT-GPT (Wu et al., 2023b). Besides, we also incorporate a concatenative approach comprising Whisper-large-v2 (Radford et al., 2023) and GPT-4 Turbo (OpenAI, 2023) for tasks related to speech as a baseline. We evaluate the performance of all these models on both fundamental and chat benchmarks, utilizing their latest publicly available checkpoints. In cases of multiple checkpoints, we select the model with the largest parameter size. For all models, we directly follow their default decoding strategies for evaluation.
   (a) Human evaluation for                  (b) Human evaluation for                  (c) Positional bias of
@@
Figure 4: The experiments of human evaluation and the position bias of GPT-4 evaluator. Figure (a) and (b) are the results of consistency between the GPT-4 evaluator and human judgement on the foundation benchmark and chat benchmark, respectively. Figure (c) refers to the result of scores by interchanging the position of the hypothesis and reference during evaluation on the chat benchmark.
@@ 4.2 Main Results
The results of LALMs are present in Table 3. For the foundation benchmark, we also conduct a comparison between the use of an exact matching strategy with our proposed GPT-4 alignment strategy. As an example, we try to match ‘B’, ‘B.’, ‘B)’, etc. with LALMs’ hypothesis for the exact matching. The results are shown in Table 4. We can find that BLSP and SALMONN have a high success rate in directly generating the choice, showcasing their strong ability to follow single-choice instruction. However, we find that it is challenging to precisely extract the predicted choice from the hypotheses of other models due to significant variations in the output formats of different LALMs. However, with the assistance of GPT-4 as the evaluator, the success rate for all models can be improved to 100%.
@@
Based on these results, we have several observations: 1) The existing LALMs either have limited audio understanding or instruction-following capabilities. For instance, Qwen-Audio Turbo achieves the highest average score in both foundation and chat benchmarks while the model displays a weak proficiency in following single-choice instructions such as often directly generating a full sentence semantically akin to one of the choices, and thus receives a low success rate for the exact matching; 2) As for chat abilities related only to speech transcription, none of the models surpass the concatenative baseline Whisper plus GPT-4.
@@ 4.3 Human Evaluation
To evaluate the consistency between the evaluations of GPT-4 and human judgements, we design experiments for both the foundation and chat benchmarks. For the foundation benchmark, we instruct the testers to determine which option aligns closest with the hypothesis. We then compare the option selected by human testers with the option chosen by GPT-4 to assess the extent of agreement. For this consistency analysis, we employed Qwen-Audio-Chat as a representative model and randomly selected 400 questions from the benchmark. These questions were then evaluated by three native English speakers. Additionally, we also compared the performance of GPT-4 with GPT-3.5 Turbo. As depicted in Figure 4 (a), GPT-4 Turbo, serving as the evaluator, exhibited a high level of consistency at 98.2% with human judgements. Comparatively, GPT-3.5 Turbo had a slightly lower consistency rate of 96.4%.
Regarding the chat benchmark, obtaining a numerical score on a scale of 1 to 10 directly from testers poses challenges. Therefore, we resort to a pairwise comparison of the models instead. Testers listen to audio and compare the performance of both models based on their usefulness, relevance, accuracy and comprehensiveness to the given question, indicating their preference as either “A is better”, “B is better”, or “Both are equal”. Subsequently, we convert the GPT-4 scores into the same preference-based rating as the human testers for any two models. We then assess the consistency between the two sets of results. For the chat benchmark, we conduct pairwise comparisons among Qwen-Audio-Chat, SALMONN, BLSP, and GPT-4. We randomly select 200 questions and have them evaluated by three native English speakers. As depicted in Figure 4 (b), the pairwise preference consistency scored above 70%, demonstrating a high level of agreement.

### v2

@@ 4 Experiments
Table 4: The success rate of different strategies of matching hypotheses with the golden choices for the foundation benchmark. The success rate denotes the probability that the model successfully responds to one of the choices.
@@ 4.1 Models
We evaluate the performance of various LALMs with instruction-following capabilities. These models are either open-sourced or accessible through public APIs, such as SpeechGPT (Zhang et al., 2023), BLSP (Wang et al., 2023a), SALMONN (Tang et al., 2023a), Qwen-Audio-Chat (Chu et al., 2023), and Qwen-Audio Turbo  [^3] https://help.aliyun.com/zh/dashscope/developer-reference/qwen-audio-api. Additionally, we consider large multi-modality models with audio understanding abilities like PandaGPT (Su et al., 2023), Macaw-LLM (Lyu et al., 2023), and NExT-GPT (Wu et al., 2023b). Besides, we also incorporate a sequential approach comprising Whisper-large-v2 (Radford et al., 2023) and GPT-4 Turbo (OpenAI, 2023) for tasks related to speech as a baseline. We evaluate the performance of all these models on both fundamental and chat benchmarks, utilizing their latest publicly available checkpoints. In cases of multiple checkpoints, we select the model with the largest parameter size. For all models, we directly follow their default decoding strategies for evaluation.
     (a)                    (b) Human evaluation for                    (c) Positional bias of
@@
Figure 4: The experiments of human evaluation and the position bias of GPT-4 evaluator. Figure (a) and (b) are the results of consistency between the GPT-4 evaluator and human judgment on the foundation benchmark and chat benchmark, respectively. Figure (c) refers to the result of scores by interchanging the position of the hypothesis and reference during evaluation on the chat benchmark.
@@ 4.2 Main Results
The results of LALMs are presented in Table 3. The detailed results are shown in Table 6. For the foundation benchmark, we also conduct a comparison between the use of an exact matching strategy with our proposed GPT-4 alignment strategy. As an example, we try to match ‘B’, ‘B.’, ‘B)’, etc. with LALMs’ hypothesis for the exact matching. The results are shown in Table 4. We can find that BLSP and SALMONN have a high success rate in directly generating the choice, showcasing their strong ability to follow single-choice instruction. However, we find that it is challenging to precisely extract the predicted choice from the hypotheses of other models due to significant variations in the output formats of different LALMs. However, with the assistance of GPT-4 as the evaluator, the success rate for all models can be improved to 100%.
@@
Based on these results, we have several observations: 1) The existing LALMs either have limited audio understanding or instruction-following capabilities. For instance, Qwen-Audio Turbo achieves the highest average score in both foundation and chat benchmarks while the model displays a weak proficiency in following single-choice instructions such as often directly generating a full sentence semantically akin to one of the choices, and thus receives a low success rate for the exact matching; 2) As for chat abilities related only to speech transcription, none of the models surpass the sequential baseline Whisper plus GPT-4.
@@ 4.3 Human Evaluation
To evaluate the consistency between the evaluations of GPT-4 and human judgments, we design experiments for both the foundation and chat benchmarks. For the foundation benchmark, we instruct the testers to determine which option aligns closest with the hypothesis. We then compare the option selected by human testers with the option chosen by GPT-4 to assess the extent of agreement. For this consistency analysis, we employed Qwen-Audio-Chat as a representative model and randomly selected 400 questions from the benchmark. These questions were then evaluated by three native English speakers. Additionally, we also compared the performance of GPT-4 with GPT-3.5 Turbo. As depicted in Figure 4 (a), GPT-4 Turbo, serving as the evaluator, exhibited a high level of consistency at 98.2% with human judgments. Comparatively, GPT-3.5 Turbo had a slightly lower consistency rate of 96.4%.
Regarding the chat benchmark, obtaining a numerical score on a scale of 1 to 10 directly from testers poses challenges. Therefore, we resort to a pairwise comparison of the models instead. Testers listen to audio and compare the performance of both models based on their usefulness, relevance, accuracy, and comprehensiveness to the given question, indicating their preference as either “A is better”, “B is better”, or “Both are equal”. Subsequently, we convert the GPT-4 scores into the same preference-based rating as the human testers for any two models. We then assess the consistency between the two sets of results. For the chat benchmark, we conduct pairwise comparisons among Qwen-Audio-Chat, SALMONN, BLSP, and GPT-4. We randomly select 200 questions and have them evaluated by three native English speakers. As depicted in Figure 4 (b), the pairwise preference consistency scored above 70%, demonstrating a high level of agreement.
@@ 5 Conclusion
6 Limitations
The objective of AIR-Bench is to develop a large-scale, extensive and generative evaluation framework that encompasses a wide range of audio domains and tasks. However, AIR-Bench currently has several limitations. Firstly, it does not incorporate tasks involving multiple audio comparisons, such as assessing music coherence, for both the foundation and chat benchmark. Besides, AIR-Bench does not encompass the evaluation of multi-turn dialogues that may involve multiple audio inputs. For evaluation, AIR-Bench relies on a powerful and robust evaluator such as GPT-4. However, the availability and accessibility of the GPT-4 API are external factors beyond our control. In the event that GPT-4 transitions to a closed-source model or implements higher pricing standards in the future, alternative evaluators will need to be explored and considered.
7 Ethical Considerations
The AIR-Bench initiative uses publicly available datasets to create a collection of relevant question-and-answer data. It then uses automated methods to evaluate this data, which is a more efficient alternative to manually evaluating it. However, there are challenges with this automated evaluation approach, including the potential for data misuse and the introduction of biases. To prevent data misuse, we follow the licenses and usage guidelines associated with the original open-source materials when generating related data. It’s important to point out that the automated evaluation could be biased. These biases may come from the datasets themselves or the scoring algorithms used, causing differences between automated evaluation results and human judgment. Therefore, the outcomes obtained from automated evaluations should be viewed with caution and used as a general benchmark, rather than a definitive measure.
@@ Appendix A Detailed Results of Foundation Benchmark
Appendix B GPT Prompts for the Chat benchmark
In Figure 5, we display the carefully crafted prompts that we have developed on our chat benchmark. The figure is divided into two sections, the upper section contains prompts designed specifically for generating question-answer pairs related to reasoning, while the lower section features prompts aimed at assessing the chat performance scores of the models.
When generating questions and reference answers, we guide the process by specifying the type of questions to be elicited, allowing GPT-4 to automatically exclude data that is less amenable to question formulation. For the evaluation of the chat performance scores, we instruct GPT-4 to take a multifaceted approach, scoring both the reference answers and the model responses. This ensures that the reference answers consistently serve as a standard for comparison.
Appendix C Prompts Engineering for GPT Scoring
In this section, we partially demonstrate the process of adjusting the prompt aimed at assessing the chat performance scores of the models.
  • If we streamline our prompt by removing the descriptions pertaining to helpfulness, relevance, accuracy, and comprehensiveness, specifically by omitting "Please rate the helpfulness, relevance, accuracy, and comprehensiveness of their responses." and "In the subsequent line, please provide a comprehensive explanation of your evaluation, avoiding any potential bias and ensuring that the order in which the responses were presented does not affect your judgment.", we found that across multiple tests, many responses that were originally scored a perfect 10 were downgraded to a 9 or 8, while the unequivocally incorrect responses saw their scores rise from an initial 1 to a 2 or 3. This suggests that including these ’superfluous’ descriptions aids the model in assigning more precise scores during the evaluation process and helps to avoid ’normalization’ of scores.
  • If we change the positioning, such as moving the entire [Detailed Audio Description] section behind the [Question] and [Answer], or swapping the positions of [Question] and [Answer]; these alterations impact the scoring, turning originally correct evaluations incorrect. Absolutely correct answers were inexplicably awarded scores as low as 5, whereas absolutely incorrect responses occasionally received scores around 5 as well. Therefore, our conclusion is that the prompt exhibits a strong sensitivity to the permutation of positions. Minor punctuation or grammatical errors do not affect the scoring.
Appendix D Examples of the Foundation Benchmark
In Table 5, we present data examples for each task within the foundation benchmark.
Appendix E Examples of LALMs’ responses
In Figure 6, we illustrate a representative response from various models on the foundation benchmark. The upper portion of the figure displays the question along with the metadata for the corresponding audio. This metadata is not provided as input to the models under evaluation, the models only have access to the audio and the question posed. The lower two columns of the figure document the responses from the 9 models being tested. Similarly, an example of responses from various models on the chat benchmark can be seen in Figure 7.
Appendix F Details in Human Evaluation
We conducted a pairwise crowd worker evaluation to assess the alignment between the judgments derived from GPT-4 and those of human evaluators for both the foundation and chat benchmarks. Each pair of evaluations was scrutinized by three native English-speaking judges. During the evaluation process, we required that the entire test be conducted in a quiet environment, with human evaluators wearing headphones to listen to the audio and to isolate noise. After obtaining the test results, we conducted sample feedback; if we identified any instances of erroneous annotations, we would report back to the outsourcing platform for them to carry out a re-evaluation.
  • For the foundation benchmark, we randomly selected 400 questions from the pool of model responses. These were accompanied by both GPT-3.5 and GPT-4 alignment results. Evaluators were instructed to ascertain whether the responses provided by GPT-3.5 Turbo and GPT-4 Turbo was accurate. The screenshots of instructions for the foundation benchmark is shown in Figure 8.
  • For the chat benchmark, we randomly chose 200 dialogues from the responses generated by Qwen-Audio-Chat, SALMONN, BLSP, and GPT-4, respectively. Evaluators were tasked with determining which model exhibited superior or equivalent performance. The screenshots of instructions for the chat benchmark is shown in Figure 9.
  • For the chat benchmark, we further analyzed correlation with human judgment based on task and audio type. After conducting a statistical analysis of the randomly selected QA pairs, we found that Speech accounts for 42%, Sound for 22%, Music for 16%, and Mixed Audio for 20%. To further confirm the association between human judgment and audio type, we categorized the results from Figure 4(b) by audio type. As shown in Table 7, the statistical results presented in the table indicate that QAs involving Music and Mixed Audio categories tend to have slightly higher alignment most of the time, whereas QAs involving Sound and Speech categories tend to have slightly lower alignment most of the time. We speculate that the reasons for the discrepancies might be: there are many situational questions in the Sound category QAs (such as ’What would you do if you heard this sound’), and many reasoning questions in the Speech category QAs. These more complex questions pose relatively greater challenges for GPT’s evaluation.
Types Task Question Example Choice Example
Speech Speech Grounding Choose when ‘hate’ is spoken.
A.[7.67, 8.05] B.[1.03, 1.53]
C.[3.07, 3.27] D.[7.02, 7.21]
Spoken language identification Recognize the language of the speech. A.en B.ja C.de D.fr
Speaker gender recognition
Detect the gender of the speaker in this audio file. A.male B.female
Emotion recognition What emotion is at the forefront of the speaker’s words?
A.angry B.happy
C.sad D.neutral
Speaker age prediction Which age range do you believe best matches the speaker’s voice?
A.teens to twenties
B.thirties to forties
C.fifties to sixties
D.seventies to eighties
Speech entity recognition Tell me the first ‘transport_type’-connected word in this audio.
A.go B.how
C.metro D.train
Intent classification What’s your opinion on the speaker’s goal in this sound clip?
A.audio_volume_up
B.news_query
C.lists_createoradd
D.play_podcasts
Speaker number verification The speech features how many speakers? A.2 B.4 C.3 D.1
Synthesized voice detection Based on your assessment, is this speech Real or Fake? A.fake B.real
Sound Audio grounding
What are the exact times when ‘a woman briefly talks’ is
present in the clip?
A.[0.44, 2.38]
B. [3.85, 4.11]
C. [9.01, 10.02]
D. [4.15, 7.83]
Vocal sound classification What’s the provenance of the sound in this clip?
A.Sigh B.Throat clearing
C.Cough D.Sneeze
Acoustic scene classification What venue are the sounds indicative of?
A.kitchen B.elevator
C.street D.crowded indoor
Sound question answering What animal makes a sound in the video?
A.cattle B.horse
C.cat D.bird
Music Music instruments classification Discern the principal instrument in this tune.
A.bass B.string
C.brass D.mallet
Music genre classification What’s the genre identity of this music?
A.Jazz B.Rock
C.Country D.Experimental
Music note analysis-pitch What is the MIDI pitch level of the note played?
A.midi_pitch_19
B.midi_pitch_29
C.midi_pitch_37
D.midi_pitch_71
Music note analysis-velocity What numerical value is the MIDI velocity for this note?
A.midi_velocity_127
B.midi_velocity_50
C.midi_velocity_100
D.midi_velocity_25
Music question answering Is the guzheng louder than the piano? A.yes B.no C.four D.one
Music emotion detection What kind of sentiment does this music invoke?
A.meditative B.positive
C.trailer D.advertising
Table 5: Examples of questions and choices on the foundation benchmark.
@@ Appendix F Details in Human Evaluation
Figure 5: GPT prompts for creating QA in the foundation benchmark and scoring in the chat benchmark.
Figure 6: The illustration of the models’ responses on the foundation benchmark.
Figure 7: The illustration of the model’s responses on the chat benchmark.
Type
GPT-4 vs
BLSP
GPT-4 vs
Qw.Chat
GPT-4 vs
SALMONN
SALMONN
vs BLSP
SALMONN
vs Qw.Chat
Qw.Chat
vs BLSP
Speech 77% 76% 89% 73% 75% 69%
Sound 73% 66% 96% 66% 75% 73%
Music 75% 88% 88% 81% 84% 75%
Mixed Audio 83% 88% 93% 75% 78% 70%
Table 7: Association between human judgment and audio type.
Figure 8: Screenshot of human evaluation for the foundation benchmark.
Figure 9: Screenshot of human evaluation for the chat benchmark.